### v2

# Heavy Neutral Leptons without Prejudice

Abstract
Heavy Neutral Leptons (HNLs) provide a compelling extension to the Standard Model, addressing the neutrino masses, baryogenesis, and dark matter problems. We perform a model-independent collider study, decoupling the active-sterile mixing angle ($V$) from the Yukawa coupling ($y$), and explore sensitivities at the HL-LHC for prompt and displaced decays. We also consider the possibility of HNLs being long-lived particles decaying in far detectors as FASER. In addition, we study the expected reach at FCC-ee for the prompt and displaced cases. For zero mixing, FCC-ee and HL-LHC sensitivities to $y$ are comparable, with Higgs width measurements imposing the strongest constraints. With non-zero mixing, sensitivities are dominated by $V$, significantly constraining parameter space. This work highlights the importance of precision Higgs studies and displaced searches in probing HNLs at current and future colliders.
1 Introduction
The discovery of the Higgs boson in 2012 [1, 2] provided significant support for the Standard Model (SM), enhancing its alignment with the observed data. Among the many triumphs of the SM are its predictions of gluons [3], the $W^{\pm}$ and $Z$ bosons’ existence and masses [4], and the identification of charm and top quarks [5, 6, 7]. However, some critical questions remain unresolved, including the origins of neutrino masses, dark matter, matter-antimatter asymmetry, and a few experimental anomalies. To address these gaps, various beyond-SM (BSM) models have been proposed, with precision Higgs physics offering potential avenues for constraint, leveraging both the HL-LHC and future Higgs factories.
Neutrino oscillation experiments have confirmed that neutrinos have mass, raising questions about potential connections to the Higgs mechanism, which is responsible for mass generation of the other SM particles. One promising BSM extension involves the addition of Heavy Neutral Leptons (HNLs), which interact with active SM neutrinos and the Higgs boson through Yukawa couplings. Depending on the model, these sterile HNLs can mix with SM neutrinos post-electroweak symmetry breaking when the Higgs acquires a vacuum expectation value (vev) [8, 9].
Several mechanisms, such as seesaw models and scotogenic approaches, can generate nonzero neutrino masses [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Beyond neutrino mass, HNLs can also contribute to the generation of the baryon asymmetry of the universe via leptogenesis [20, 21, 22, 23, 24, 25] and may even explain the observed dark matter relic abundance [26, 27, 28, 29]. In particular, introducing three HNLs into a minimal neutrino SM framework could address all these challenges [30, 22].
Instead of focusing on a specific model, this study adopts a model-independent approach to experimentally test HNLs. Many searches have probed high-mass HNLs through their production or prompt decays in colliders [31, 32, 33, 34, 35, 36, 37], see also Refs. [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. In certain parameter regions, HNLs may have substantially displaced decays within detectors, creating a unique displaced vertex signature [60]. Recent studies suggest that these signatures could be targeted in LHC searches with associated charged leptons [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73], via Higgs decays [74, 75, 76, 77, 78, 79, 80, 81], and even in detectors such as LHCb [82]. Similar efforts are underway at DUNE [83, 84], IceCube [85], proposed lepton colliders [86, 87], and SHiP [88, 89]; the latter shows particular promise for HNLs below the $c$-quark mass through meson decays [90]. ATLAS and CMS have already performed some of these analyzes [91, 92, 93, 94, 95].
Additional far detectors at the LHC, such as FASER [96, 97], MoEDAL-MAPP [98, 99], MATHUSLA [100, 101, 102], ANUBIS [103], CODEX-b [104], and FACET [105], have been proposed to further study long-lived particles, with FASER and MoEDAL-MAPP currently operational. The case of HNLs as long-lived particles has been a focus in recent studies [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118].
Future colliders could significantly enhance the sensitivity to HNLs and related phenomena compared to current facilities. The Future Circular Collider in its electron-positron mode (FCC-ee) [119] stands out as a promising candidate, due to its ability to achieve high-statistics data across various operational phases and its exceptionally clean experimental environment. This makes FCC-ee particularly well suited for probing both prompt and displaced decay signatures of HNLs.
To date, HNL searches have mainly focused on signatures from mixing with active neutrinos in the SM. However, the reliance on mixing for the HNL phenomenology is highly dependent on the chosen model. In scenarios where Yukawa couplings are dominant, Yukawa interactions could solely characterize the phenomenology of HNL. Thus, we propose an alternative approach that leverages HNL Yukawa interactions with the Higgs boson along with the mixing contribution from gauge bosons. We keep Yukawa couplings and mixing angles independent to minimize model assumptions. The prospects of the setup at the LHC treating mixing angles as negligible and the role of precision Higgs constraints were already studied in Ref. [120]; here, we include the extension of the study to FCC-ee. The aim of the current study presented here is more general, and the results help us delineate the parameter space where one dominates over the other.
The paper is organized as follows. Section 2 outlines the model and analysis setup. Section 3 describes HNL production and decay along with constraints from various experimental measurements. Sections 4.1, 4.2, and 4.3 cover analyses for the LHC of HNLs that decay promptly, with displaced vertices, and in far detectors, respectively. Section 5 deals with the prompt and displaced analyzes at FCC-ee for two key operational phases, $Z$ pole in Section 5.1, and $Zh$ production in Section 5.2. Finally, Section 6 presents our conclusions. In the appendix, we present, where relevant, the recast of existing constraints on mixing in our setup.
2 Set-up
In this work, we extend the SM by introducing $n$ HNLs, labeled $\widetilde{N}$, which can be Dirac or Majorana particles. The relevant Yukawa interaction is given by
$y_{N\alpha}\,\overline{\widetilde{N}}\,\widetilde{H}^{\dagger}\,L_{\alpha}\,,$ (2.1)
where $H$ is the SM Higgs doublet, $\widetilde{H}\equiv i\sigma^{2}H^{*}$ (with $\sigma^{2}$ as the second Pauli matrix), $L_{\alpha}$ are the SM lepton doublets for $\alpha=e,\mu,\tau$, and $y_{N\alpha}$ are the Yukawa couplings linking HNLs with SM leptons. The mass of each HNL, $m_{N}$, is introduced, but is not explicitly written here, since it depends on whether $\widetilde{N}$ is Dirac or Majorana.
After electroweak symmetry breaking, the Higgs field acquires a vev $v_{h}\simeq 246$ GeV, yielding Dirac-type mass terms for the HNLs, defined as
$(M_{D})_{N\alpha}\equiv\frac{y_{N\alpha}\,v_{h}}{\sqrt{2}}\,.$ (2.2)
Similar terms can also arise if additional scalar fields, which gain vevs, are present in the model. The physical mass eigenstates of the three SM neutrinos $\nu$ and the $n$ HNLs are then obtained by diagonalizing the full mass matrix in a manner dictated by the model. To satisfy the neutrino oscillation data, at least two HNLs ($n\geq 2$) are necessary [121, 122]. Additionally, fitting sub-eV neutrino masses [123] imposes specific constraints on the model’s parameter space. Diagonalizing to the mass basis introduces mixings between HNLs and SM neutrinos, which, if present, allows HNLs to interact with other SM particles, especially the $W^{\pm}$ and $Z$ bosons. For the type-I seesaw mechanism [10, 11, 12, 13, 14, 15, 16], the mixing angle scales with $y_{N\alpha}\,v_{h}/m_{N}$. However, in inverse seesaw and double-seesaw models [124, 125], an additional mass scale reduces the mixing contributions. In particular, some models can fine-tune mixing contributions to be precisely zero [126, 127, 120]. In contrast, in scotogenic models of neutrino mass, it is the Yukawa coupling to the SM Higgs that vanishes [128, 129]. It is important to note that while some models connect Yukawa couplings and mixing angles, these relationships should not be universally assumed. This had also been emphasized strongly in our previous study [120].
To maintain a model-independent approach, in the mass basis, we treat the Yukawa couplings, the mixing angles, and the physical HNL masses as separate free parameters, being agnostic about their possible interdependence. This approach allows for a broader exploration of the HNL parameter space. Such a set-up can be particularly relevant for displaced decay searches and precision Higgs studies.
3 Production and Decays of HNLs
HNLs with masses in the GeV range can be produced via decays of gauge and Higgs bosons. For the production via Higgs, the partial decay width of the Higgs boson into an HNL and an active neutrino is given by
$\Gamma(h\to N\nu)=\frac{y^{2}}{8\pi}\,m_{h}\left[1-\left(\frac{m_{N}}{m_{h}}\right)^{2}\right]^{2},$ (3.1)
where $m_{h}\simeq 125$ GeV is the Higgs mass, and $y^{2}\equiv y_{Ne}^{2}+y_{N\mu}^{2}+y_{N\tau}^{2}$ represents the sum of all Yukawa couplings. Since LHC is insensitive to neutrino flavor, $y^{2}$ covers all contributions. This decay mode affects the total decay width of the Higgs boson, which is predicted in the SM to be $\Gamma_{h}\simeq 4.1$ MeV [130]. Current measurements at the LHC place the Higgs width at $\Gamma_{h}=3.2^{+2.4}_{-1.7}~{}\text{MeV}$ [131]. The high-luminosity LHC (HL-LHC) is projected to significantly improve this precision, achieving an uncertainty of just $5.3\%$ [132]. It is expected to reduce further to $1\%$ at FCC-ee [133]. In addition, for the $W$- and $Z$-boson decays, the partial widths are
$\Gamma(W^{\pm}\to Nl_{\alpha}^{\pm})$ $\simeq\frac{e^{2}\,V_{N\alpha}^{2}}{96\pi\,s_{W}^{2}}\,m_{W}\left[2-3\left(\frac{m_{N}}{m_{W}}\right)^{2}+\left(\frac{m_{N}}{m_{W}}\right)^{6}\right],$ (3.2)
$\Gamma(Z\to N\nu)$ $=\frac{e^{2}\,V^{2}}{96\pi\,c_{W}^{2}\,s_{W}^{2}}\,m_{Z}\left[2-3\left(\frac{m_{N}}{m_{Z}}\right)^{2}+\left(\frac{m_{N}}{m_{Z}}\right)^{6}\right],$ (3.3)
where $V^{2}\equiv V_{Ne}^{2}+V_{N\mu}^{2}+V_{N\tau}^{2}$. The SM total decay widths for the $Z$ and $W$ bosons are $\Gamma_{Z}=(2.4940\pm 0.0009)$ GeV and $\Gamma_{W}=(2.0892\pm 0.0008)$ GeV, respectively, while their decay into SM neutrinos has partial widths $\Gamma(Z\to\nu\bar{\nu})=(501\pm 0.045)$ MeV and $\Gamma(W^{\pm}\to\nu_{\alpha}l_{\alpha}^{\pm})=679\pm 0.12$ MeV [131]. The measured values are pretty close with $\Gamma_{Z}=(2.4955\pm 0.0023)$ GeV, $\Gamma_{W}=(2.085\pm 0.042)$ GeV, $\Gamma(Z\to\texttt{inv})=(500\pm 1.5)$ MeV and $\Gamma(W^{\pm}\to\nu_{\alpha}l_{\alpha}^{\pm})=679\pm 0.01$ MeV [131].
Figure 1: Contour for $c\tau=1$ mm, 1 m and 480 m, for $m_{N}=$ 2, 5, 10 and 20 GeV, in the plane $[y^{2},V^{2}]$. The blue dotted line corresponds to the type-I seesaw relation. The red dotted line represents the contour for equal decay width through the Yukawa and mixing. The red bands are in tension with the experimental measurements. The bounds on $V^{2}$ are taken from the most stringent limits on mixing [134, 135]. The bounds on $y^{2}$ arise from the total and invisible decay width of the Higgs.
Once produced, HNLs in the GeV ballpark and below decay into three fermions via off-shell gauge or Higgs bosons. We compute these decay widths in MADGRAPH5_aMC@NLO v2.9.16 [136, 137] using a modified version of the FeynRules [138, 139] model HeavyN [41, 45, 50]. Higgs-mediated decays favor heavy fermions (e.g. $b\bar{b}$, $\tau^{+}\tau^{-}$, $c\bar{c}$) for $m_{N}$ above their thresholds, while gauge boson-mediated decays are universal in this respect. Figure 1 shows the contours for the proper length $c\,\tau=1$ mm, 1 m, and 480 m, for HNLs in the plane $[y^{2},V^{2}]$, corresponding to the prompt, displaced and long-lived regimes. Increasing $m_{N}$ reduces $c\tau\propto 1/m_{N}^{5}$, shifting the contours toward lower coupling values.
The vertical bounds for $y^{2}$, shown in red, arise from the Higgs total and invisible decay width constraints. Constraints from the invisible branching fraction of the Higgs arise only when the HNL decays through the Yukawa coupling outside the detector [120]. ATLAS and CMS limits on $\text{Br}(h\to\text{inv})$ are $<10.7\%$ and $<15\%$, respectively, at 95% CL [140, 141], providing stronger constraints than the Higgs total width for $m_{N}\lesssim 10$ GeV, and are represented by the rectangular vertical red region with rounded edges. These limits are expected to improve to a precision of $1.9\%$ at HL-LHC and $0.2\%$ at FCC-ee [142]. Invisible decays of HNLs [41] can also arise through mixing
$\Gamma(N\to\nu\nu\nu)=\frac{G_{F}^{2}}{96\pi^{3}}V^{2}m_{N}^{5},$ (3.4)
but it contributes minimally (approximately 6%) [70]. These decays can be constrained by the invisible width of the $Z$ boson [143] to derive limits on $V^{2}$. However, the most stringent limits on $V^{2}$ (shown by the horizontal red region) stem from current experimental bounds at colliders [134, 135]. The blue dotted line represents the type-I seesaw relation ($V=y\,v_{h}/m_{N}$) as a reference. The red dotted line corresponds to the equal decay width through the mixing and the Yukawa.
4 Prospects at LHC
In the following, projections for the HL-LHC will be presented in the case of prompt decays, displaced decays inside the ATLAS or CMS detectors, and long-lived particles.
4.1 Prompt Decays
The production of single-Higgs bosons in the LHC is dominated by gluon fusion (ggF) and vector boson fusion (VBF), with corresponding cross sections $\sigma_{\text{ggF}}\simeq 54.8$ pb and $\sigma_{\text{VBF}}\simeq 4.26$ pb at $\sqrt{s}=14$ TeV [144, 145]. The unique topology of VBF, featuring two forward jets ($j_{1},j_{2}$) in opposite hemispheres, large pseudo-rapidity separation ($\Delta\eta_{j_{1}j_{2}}$), and high invariant mass ($m_{j_{1}j_{2}}$), allows a large suppression of QCD multi-jet backgrounds. We explore single-Higgs boson production via VBF at an integrated luminosity of $\mathcal{L}=3$ ab-1. Subsequently, the Higgs can decay into an HNL and a SM neutrino, with the HNL later decaying into $b\bar{b}$ and another active neutrino via an off-shell Higgs or gauge boson (which can only be $Z$ for the final state of our interest).
The dominant backgrounds include the production of Higgs through VBF and ggF with a later decay $h\to b\bar{b}$, $t\bar{t}$ + jets, and $b\bar{b}$ + jets. The ggF background can mimic the VBF via hard forward jets, despite a distinct topology. The $t\bar{t}$ + jets process is significant due to its large cross section (990 pb at NNLO [146]) and the inherent $b$ jets. Similarly, $b\bar{b}$ + jets, with a cross section of $\sim 10^{5}$ pb [147], can mimic the VBF topology in specific configurations.
To distinguish the signal from the background, we impose the following selection criteria [120]:
  • S1: Events must have no charged lepton or photon candidates.
  • S2: At least two $b$-tagged jets and two non-$b$ jets are required.
  • S3: Leading non-$b$ jets must satisfy $p_{T_{j_{1}}}>60$ GeV, $p_{T_{j_{2}}}>40$ GeV, and $H_{T}>140$ GeV for all non-$b$ jets.
  • S4: The two leading non-$b$ jets must meet VBF criteria: $\eta_{j_{1}}\times\eta_{j_{2}}<0$, $\Delta\eta_{j_{1}j_{2}}>3.5$, $m_{j_{1}j_{2}}>500$ GeV, and $\Delta\phi_{j_{1}j_{2}}<2.5$.
  • S5: Additional non-$b$ jets ($j_{i}$) are tested using the ratio:
    $m^{\text{rel}}_{j_{i}}\equiv\frac{\text{min}(m_{j_{1}j_{i}},\,m_{j_{2}j_{i}})}{m_{j_{1}j_{2}}}\,,$ (4.1)
    where smaller values of $m^{\text{rel}}_{j_{i}}$ ($m^{\text{rel}}_{j_{i}}<0.08$) indicate compatibility with final-state radiation, helping the rejection of QCD multijet events.
[…]
  • S7: The invariant mass of the $b$ jets must satisfy $0.2\,m_{N}\leq m_{b\bar{b}}\leq 1.6\,m_{N}$, with $\Delta R_{b\bar{b}}\leq 2.5$.
Figure 2: Discovery reach at $3~{}\sigma$ CL for HL-LHC with VBF Higgs with both Yukawa coupling and mixing angle HNL of mass 45 GeV (left) and 70 GeV (right). The regions to the right of the $3\sigma$ contours will be probed at HL-LHC. The horizontal and vertical red bands arise from the current most-stringent limit on mixing and Higgs total decay width constraint on the Yukawa respectively.
Figure 2 presents the discovery reach for the HL-LHC, for two benchmark values of $m_{N}$, 45 GeV (left) and 70 GeV (right), and the selection cuts previously mentioned. HL-LHC will be able to probe the regions to the right of the $3~{}\sigma$ contours. Since HNLs are produced via the Higgs boson, achieving a $3\sigma$ significance requires a minimum $y^{2}$, resulting in the near-vertical contour. As $V^{2}$ increases, mixing-mediated decays become more efficient, leading to a reduction of the branching fraction of the Higgs into a couple of $b$ quarks, and therefore to a change in the slope of the contour. The red bands show the parameter regions that are in tension with the current experimental limits on mixing and the Higgs total decay width on the Yukawa.
4.2 Displaced Vertices
This section addresses the scenario in which HNLs decay within the inner tracker of the ATLAS or CMS detectors, resulting in a displaced vertex. As before, we concentrate on Higgs production via VBF to identify the distinct signature of two prominent forward jets. However, ggF events can also significantly meet the same topology criteria, and thus they have been incorporated into the analysis. The selection criteria for the events are as follows:
  • Same cuts $S_{1}$, $S_{3}$ and $S_{4}$ for the VBF hard forward jets, as described in Section 4.1.
  • Two extra jets with a displacement of 1 mm $\leq d_{xy}\leq 1$ m and $d_{z}\leq 300$ mm.
Initially, we consider a scenario without background, utilizing a Poisson distribution. We then focus on the parameter space where there are more than 3.09 anticipated signal events at 95% CL [131]. Figure 3 shows the region within a thick solid black line, the sensitivity reach of the searches for a displaced decay for $\sqrt{s}=14$ TeV and $\mathcal{L}=3$ ab-1, in the plane $[y^{2},\,V^{2}]$, for $m_{N}=10$ GeV (left) and $m_{N}=20$ GeV (right). These regions show the same trend as already pointed out in Fig. 1, but smaller values of $y^{2}$ are not allowed due to the insufficient number of HNLs produced.
Figure 3: Sensitivity reach of searches for a displaced decay, for $\sqrt{s}=14$ TeV and $\mathcal{L}=3$ ab-1 are shown by regions inside the contours. The thick line corresponds to the assumption of zero background events, whereas for the thin line 65 events were assumed; see the text for further details. The blue dotted line corresponds to the type-I seesaw relation. The red bands are in tension with the experimental measurements.
The hypothesis of having no background previously considered might be overly optimistic, as background events can stem from incorrectly identified displaced events due to detector resolution issues or random track intersections. To get a general idea of the potential impact of including a background on sensitivity, we refer to the discussions in Refs. [80, 70, 71, 72] and consider the maximum possible background, consistent with the lack of observed background events in the ATLAS analysis found in Refs. [148, 149]. This corresponds to a maximum of 3 background events for $\mathcal{L}=139$ fb-1, which scales to 65 background events for a luminosity of 3 ab-1. This pessimistic case is illustrated in Fig. 3 with a thin black line.
4.3 Long-lived Particles
The FASER detector is located 480 meters downstream of the proton-proton interaction point leveraged by the ATLAS experiment [96, 97]. Designed to detect new particles within a cylindrical region of radius 10 cm and length 1.5 m, it will initially function at an integrated luminosity of 150 fb-1. In contrast, the second phase (FASER-2 [97]) at the HL-LHC will have a length of 10 m, a radius of 1 m, and an integrated luminosity of 3 ab-1. Our focus here is on this second phase.
Detection of long-lived HNLs can be performed using a detector such as FASER [96], which is adept at detecting particles decaying approximately 480 m from ATLAS interaction point. This considerable separation ensures minimal background noise, often deemed negligible. Consequently, the experiment achieves its 95% CL sensitivity with only 3.09 events [131].
In this section, we will explore HNLs generated from Higgs decay processes. The overall inclusive cross section for a single Higgs boson is mainly influenced by the ggF and VBF mechanisms, with $\sigma\simeq 59.1$ pb at $\sqrt{s}=14$ TeV [144, 145]. As before, we will consider the total integrated luminosity of $\mathcal{L}=3$ ab-1.
Given the geometry of FASER, the probability $\mathcal{P}$ of a HNL decaying inside the detector is
$\mathcal{P}=\left[e^{-(L-\Delta)/d}-e^{-L/d}\right]\,\Theta\left[R-L\,\tan\theta\right],$ (4.2)
where $d$ is the decay length in the laboratory frame of the HNL, and $\theta$ the angle between the momentum and the beam line. This decay length takes into account the Lorentz boost factor
$d=c\tau\,\beta\,\gamma=c\tau\,\frac{|\vec{p}|}{m_{N}}\,.$ (4.3)
Figure 4: The sensitivity reaches for FASER are represented by the region inside the solid black lines, for $\sqrt{s}=14$ TeV and $\mathcal{L}=3$ ab-1. The blue dotted line corresponds to the type-I seesaw relation. The red bands are in tension with the experimental measurements.
Figure 4 shows, with a thick solid black line, the sensitivity reach of FASER-2 to HNLs of masses 2, 5, 10 and 20 GeV in the plane $[y^{2},\,|V|^{2}]$, where the parameter regions inside the contours will be probed. Here, the HNLs can be produced from the decays of Higgs and gauge bosons, and we allow all possible decay modes of the HNL. For completeness, the partial contributions of the different production modes are also shown with thin lines.
5 FCC-ee
The initial phase of the FCC integrated project involves an $e^{+}e^{-}$ collider known as FCC-ee [119]. This collider is designed to function as a Higgs factory, as well as an electroweak and top factory, delivering the highest luminosities across four distinct center-of-mass energy stages: the $Z$ pole, the $W^{+}W^{-}$ threshold, the $Zh$ production peak, and the $t\bar{t}$ threshold. Our focus lies on the $Z$-pole and $Zh$-production stages to evaluate our setup. The $Z$-pole stage is particularly significant due to the enormous number of $Z$ bosons expected to be produced by the end of its operation ($\sim 6\times 10^{12}$ $Z$ bosons), providing a substantial source of HNLs, which can subsequently decay through off-shell gauge bosons and Higgs. Meanwhile, the $Zh$ production mode offers a direct mechanism to produce HNLs via Higgs decays, serving as a complementary probe to the zero-mixing scenario proposed in Ref. [120].
5.1 $Z$ Pole
The $Z$-pole phase of the FCC-ee corresponds to a center-of-mass energy of $\sqrt{s}=91.2$ GeV. This phase is characterized by exceptionally high luminosity, making it an ideal environment for precision measurements and rare process searches. During the first two years of operation, the collider is expected to achieve a luminosity of $\mathcal{L}=34$ ab-1 per year. This luminosity will double in the following two years, reaching $\mathcal{L}=68$ ab-1 per year. Therefore, the total integrated luminosity for this phase will be $\mathcal{L}=204$ ab-1.
The process of interest at the $Z$ pole is the production of HNLs through $Z$ decays. The relevant process chain can be expressed as [^1] HNLs decaying to $\mu jj$ final state (through an off-shell $W$) could have a greater reach in the $V^{2}$ vs. $m_{N}$ plane. However, we do not focus on that channel as the Yukawa coupling contribution is very suppressed.
$e^{+}e^{-}\to Z\to\nu N\to\nu\bar{\nu}b\bar{b}\,.$ (5.1)
[…]
In the next subsections, the cases where the HNL decays promptly and with a displaced vertex inside the inner tracker will be analyzed separately.
5.1.1 Prompt Decays
The optimization of the selection regions for the prompt decays includes the following cuts:
  • S1: Event must have at least two $b$-tagged jets.
  • S2: The two leading $b$ jets must satisfy $p_{T_{b_{1}}}<60$ GeV and $p_{T_{b_{2}}}<40$ GeV.
[…]
  • S4: The invariant mass of the $b$ jets must satisfy $0.2\,m_{N}\leq m_{b\bar{b}}\leq 1.5\,m_{N}$, with $\Delta R_{b\bar{b}}\leq 2.5$.
Figure 5: Discovery reach for prompt decays at FCC-ee $Z$-pole phase represented by regions above the black contours for three benchmark masses of the HNL, namely 45 and 70 GeV.
Figure 5 illustrates the results for prompt decays of HNLs corresponding to the benchmark mass values: $m_{N}=45$ and 70 GeV. The production of HNLs requires a minimum value of the mixing parameter $V^{2}$, which leads to horizontal contours in the $[V^{2},\,y^{2}]$ plane. For higher values of the Yukawa coupling $y$, the decay through the off-shell Higgs becomes dominant, as indicated by the change in the slope of the contours.
The benchmark mass point $m_{N}=45$ GeV is already excluded by current experimental constraints on the mixing parameter $V^{2}$. For $m_{N}=70$ GeV, the parameter space is still allowed within the current bounds on the mixing. However, the region of interest, where the Yukawa coupling $y$ significantly influences the HNL decay dynamics, is already ruled out by constraints on the width of the Higgs boson $\Gamma_{h}$.
5.1.2 Displaced Decays
Figure 6 shows the results of the search for displaced vertex for HNLs with masses $m_{N}=2$, 5, 10, and 20 GeV. The analysis requires two jets with transverse displacements satisfying $1~{}\mathrm{mm}\leq d_{xy}\leq 1$ m and longitudinal displacement $d_{z}\leq 300$ mm. The detection sensitivities are illustrated for 1, 10, and 100 observed events, corresponding to dot-dashed, dashed, and solid contours, respectively.
Figure 6: Sensitivity reaches for displaced vertices at FCCee in the $Z$-pole phase for all channels combined are shown by the regions inside the thick, dashed and dot-dashed contours for 100, 10 and 1 events respectively. The blue dotted line corresponds to the type-I seesaw relation. The red bands are in tension with the experimental measurements.
For $m_{N}=2$ GeV and $m_{N}=5$ GeV, the parameter space is already excluded by current experimental constraints on mixing and Higgs measurements. For $m_{N}=10$ GeV, some parameter space remains accessible for mixing-sensitive searches, but the region where the Yukawa coupling $y$ plays a significant role is excluded. The most promising case is $m_{N}=20$ GeV, where a substantial portion of the parameter space relevant to our setup is still allowed by the current experimental limits.
5.2 $Zh$ Production
The $Zh$ production mode at FCC-ee is characterized by a center-of-mass energy of $\sqrt{s}=240$ GeV, which allows for precise studies of the Higgs boson and its interactions. The planned integrated luminosity for this phase is $\mathcal{L}=2.4$ ab-1 per year for a period of three years, which leads to a total integrated luminosity of $\mathcal{L}=7.2$ ab-1.
The production cross section for the $Zh$ process is $\sigma(e^{+}e^{-}\to Zh)=0.2403$ pb, making this mode a significant channel for Higgs studies. The associated $Z$ boson, when decaying into $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$, provides a clean experimental signature for the $Zh$ process. The main background contributions come from the direct decay of $Z$ or $h$ to the charged fermion final states of the signal of which we are interested.
In the context of HNL searches, the $Zh$ production mode allows direct production of HNLs through Higgs decays, that is, $h\to N\nu$. This serves as a complementary probe to other setups, such as those based on the $Z$-pole production discussed earlier, or the zero-mixing case of HNL production through Higgs at the LHC, potentially covering regions of parameter space that are inaccessible in the $Z$-pole phase alone.
5.2.1 Without Mixing
For the zero-mixing case in the $Zh$ production mode at FCC-ee, the analysis focuses on prompt and displaced decays of the HNL. The prompt decay selection is based on the four criteria discussed in Section 5.1.1, combined with the requirement of at least two additional leptons arising from the decay of the associated $Z$ boson. For displaced decays, the selection criteria remain consistent with those outlined in Section 5.1.2, focusing on displaced vertices with measurable transverse and longitudinal displacements.
Figure 7: Left: Discovery reach comparison of Yukawa-only case between FCC-ee $Zh$ production (blue) and HL-LHC VBF Higgs production (black). Projections for the Higgs total and invisible width at HL-LHC and FCC-ee are also shown by the dot-dashed and dotted contours, respectively. Right: Comparison of the discovery reaches between HNL production at the Higgs pole and the $Zh$ prompt case, at FCC-ee.
The results of this analysis are illustrated in the left plot of Fig. 7. The blue contours represent the reach for both prompt and displaced decays. Compared to the sensitivity of HL-LHC for the zero-mixing case using VBF Higgs production [120] (represented by the black contours), the FCC-ee $Zh$ production mode does not provide an improvement in reach. The future projections on the invisible and total decay width of Higgs at FCC-ee, shown by the lower dotted and dot-dashed contours, can probe the prompt case completely along with a considerable amount of parameter space for the displaced case.
In the right panel of Fig. 7, the sensitivity to resonant Higgs production at its pole mass is depicted for two prospective FCC-ee operation scenarios, corresponding to luminosities of 10 ab-1 and 35 ab-1 (black curves). Due to the small Yukawa coupling of electrons, these sensitivities remain comparatively less significant than the one for the Higgs production through Higgstrahlung in the case of $Zh$ (blue curve).
5.2.2 With Mixing: Prompt
In the general case, the HNL can be produced from either the decay of the $Z$ boson or the Higgs boson. The analysis therefore considers both the production channels and their respective features. For HNL production through Higgs, the selection criteria remain the same as described in Section 5.2.1.
When the HNL is produced from the $Z$ boson decay, the Higgs decay into two leptons is considered, primarily via the $\tau^{+}\tau^{-}$ channel. For this case, the decay of the Higgs directly into $b\bar{b}$ is deliberately avoided to maintain consistency in the final state and to circumvent the suppression arising from the efficiency of $b$-tagging, which would scale as the fourth power or more due to the presence of a minimum of four $b$ jets.
Figure 8: Discovery reach represented by regions on the right of the black contour for prompt decays at FCC-ee $Zh$ channel.
[…]
5.2.3 With Mixing: Displaced
For the displaced case, we do not apply any specific kinematic cuts to identify whether the parent particle is a $Z$ or a Higgs boson. This generalized approach allows us to consider contributions from all possible production modes of the HNL. In the case of electron flavor interactions, an additional contribution arises from the production of $N$ and $\nu_{e}$ through the $t$-channel $W$ exchange diagram. This additional production channel enhances the sensitivity reach for the electron-flavor case. The combined sensitivity reach, accounting for all these contributions, is illustrated in the top two panels of Fig. 9 for $m_{N}$ of 10 and 20 GeV. For HNLs that interact with muon or tau flavors, the $t$-channel $W$ exchange contribution is not present. The results for these cases for the same masses are shown in the bottom two panels of Fig. 9.
Figure 9: Sensitivity reaches for displaced vertices at FCCee with $\sqrt{s}=240$ GeV for the $Zh$ production mode (all channels combined) are shown by regions inside the solid, dashed and dot-dashed contours of 100, 10 and 1 events. The upper two plots correspond to HNLs coupling to the electron flavor and the bottom two plots correspond to HNLs coupling to muon and tau flavors. The blue dotted line corresponds to the type-I seesaw relation. The red bands are in tension with the experimental measurements.
6 Conclusions
The extension of the Standard Model (SM) with heavy neutral leptons (HNLs) is motivated by several compelling theoretical and phenomenological considerations, such as dark matter, baryogenesis via leptogenesis, and neutrino masses. In general, HNLs can couple to SM particles through Yukawa couplings with the Higgs, and mixings with active neutrinos. In specific models, there is a tied relation between the Yukawa couplings and the mixing angles. However, one can also find scenarios where the Yukawa coupling to the SM Higgs is zero, or where mixings with active neutrinos vanish.
In this work, we remain skeptical about the possible relation between the Yukawa couplings and the mixing angles, tackling the problem without introducing mundane prejudices about the couplings of the HNLs. In order to explore the HNL parameter space, we performed a detailed collider phenomenology study within a model-independent framework, where no direct correlation is assumed between the active-sterile mixing angle and the HNL Yukawa coupling.
Our analysis considered the sensitivity of both prompt and displaced decay signatures at the HL-LHC, utilizing interaction point detectors as well as far detectors such as FASER. We also evaluated the sensitivity reach of the FCC-ee, focusing on its operation at the $Z$ pole center-of-mass energy and during the optimized $Zh$ production mode. For the scenario with no active-sterile mixing, we found no significant enhancement in sensitivity to the Yukawa coupling when comparing the HL-LHC and FCC-ee. Notably, the precision measurement of the Higgs boson width emerges as the strongest constraint on the HNL Yukawa coupling in this case. In scenarios allowing nonzero mixing, the sensitivity is dominated by the active-sterile mixing angle, which provides stringent constraints on the parameter space, especially at FCC-ee. In general, large regions of the parameter space will be probed in HL-LHC and next-generation colliders.
All in all, these results highlight the complementarity between collider experiments in probing HNL properties and emphasize the crucial role of precision Higgs measurements in constraining new physics beyond the SM.
Acknowledgments
NB received funding from Grant PID2023-151418NB-I00 funded by MCIU/AEI/10.13039/ 501100011033/ FEDER, UE.
Appendix A Appendix
In order to assess whether existing experimental constraints on the mixing angle $V^{2}$ can impose observable limits in the general case where $y$ is treated as an independent parameter, we recast the current limits from the $[V^{2},\,m_{N}]$ plane onto the $[V^{2},\,y^{2}]$ plane, showing contours for fixed mass values. The recast is performed by matching the predicted number of events in our framework with the observed events at the experimental facilities. Since the factors of luminosity and efficiency cancels out on both sides, the equation for a fixed value of HNL mass reduces to
$V^{2}(m_{N})\times\text{BR}^{(Z^{*}/W^{*}/h^{*})}_{N\rightarrow\nu f\bar{f}}(V^{2},y^{2},m_{N})=V^{2}_{\text{exp}}(m_{N})\times\text{BR}^{(Z^{*}/W^{*})}_{N\rightarrow\nu f\bar{f}}(m_{N})\,,$ (A.1)
where $V^{2}_{\text{exp}}$ and $\text{BR}^{(Z^{*}/W^{*})}_{N\rightarrow\nu f\bar{f}}$ represents the experimental constraint and branching ratio of HNL to the final state fermions ($f$) through mixing at $m_{N}$. $V^{2}$ represents the mixing applicable to our setup at the same mass and $\text{BR}^{(Z^{*}/W^{*}/h^{*})}_{N\rightarrow\nu f\bar{f}}$ is the branching ratio of HNL to the final state fermions ($f$) in our setup, which now depends on both $V^{2}$ and $y^{2}$ for a given $m_{N}$.
Figure 10: Recast showing contours for different values of HNL mass in $[V^{2},\,y^{2}]$ plane. On the left is DELPHI at LEP and on the right is the most stringent experimental limit on the masses concerned.
We first perform this analysis using the DELPHI constraints at LEP [31], as its focus on jet-based final states aligns well with aspects of our setup. The results, shown in the left panel of Fig. 10, indicate that there is no observable impact on $V^{2}$ with increasing $y^{2}$ for values $m_{N}$ of 1, 10, 30, and 60 GeV.
Next, we extend this exercise by considering the most stringent existing bounds on the mixing angle for these four mass points, as presented in the right panel of Fig. 10. In this case, the contours exhibit a change in slope at very high values of $y^{2}$. This is somewhat expected because the HNL production in these analyses occurs via charged- or neutral-current processes. These regions are likely already excluded by constraints from the invisible decay width and total decay width of the Higgs boson, and thus provide us with no additional insight into this general setup, underscoring the importance of the HL-LHC and FCC-ee.